# Combinatorial Optimization using Comparison Oracles

### v1

@@ Abstract
    Having resolved the query complexity, we turn to the algorithmic question. Our second result is a novel Global Subspace Learning (GSL) framework tailored for objective functions with discrete integer weights bounded by $B$. We give an algorithm to sort all feasible sets by objective value using only $O(nB\log(nB))$ queries, improving upon the $\widetilde{O}(n^{2})$ bound when $B=o(n)$. Moreover, we show that this framework can be efficiently implemented for linear matroids using algebraic techniques, yielding efficient algorithms with improved query complexity for problems such as $k$-SUM, SUBSET-SUM, and A+BA{+}B sorting.
@@
    Our third set of results give the first polynomial-time, low-query algorithms for several classic combinatorial problems. We develop such algorithms for finding minimum cuts in simple graphs, minimum weight spanning trees (and matroid bases in general), bipartite matching (and matroid intersection), and shortest $s$-$t$ paths.
Our work provides the first general query complexity bounds and efficient algorithmic results for this fundamental model, opening new directions for comparison-based optimization, and suggests many directions for future research.
@@ 1.1.2 Global Subspace Learning
Unfortunately, the approach of Theorem˜1.1 requires explicitly maintaining the set of feasible solutions, which could have size exponential in $n$; we currently do not know efficient implementations for it, even for simple problems like finding the subset of size $k$ of largest weight (a.k.a. optimizing over rank-$k$ uniform matroids) for large $k$, which can be easily solved using other approaches.
In our quest for efficient algorithms, we develop another, different low-query-complexity algorithm for the case where the weights in the objective function are integers in the range [B,B][-B,B]. This algorithm is more intuitive, and in fact, sorts all the feasible sets by their weight (again, in $\operatorname{poly}(|\mathcal{F}|,n)$ time for now):
@@
The main insight behind this result is the following: two sets $S,T\in\mathcal{F}$ with the same weight satisfy $\langle w^{*},\mathbbm{1}_{S}-\mathbbm{1}_{T}\rangle=0$, and hence the vector $(\mathbbm{1}_{S}-\mathbbm{1}_{T})$ gives us information about the subspace orthogonal to $w^{*}$. Our algorithm maintains a subspace $\mathcal{A}$ spanned by these orthogonal directions $(\mathbbm{1}_{S}-\mathbbm{1}_{T})$, and uses it to infer the weight class of other sets in $\mathcal{F}$ without making any additional comparisons; the details appear in Section˜3. While Theorem˜1.3 relies on weights being integral and bounded, it highlights how the structure of linear optimization can be exploited.
This explicit linear-algebraic structure allows us to get efficient algorithms from Theorem˜1.3:
@@
Moreover, we get new algorithms for the Subset Sum, $k$-Sum and Sorting $P+Q$ problems studied by [KLM19], with better bounds for the case when $B=o(n)$; see Section˜3.3 for details.
@@ 1.1.3 Efficient algorithms for Classical Combinatorial Families
Finally, we show that cut comparisons suffice to construct sparsifiers for a broad class of graph problems called heavy subgraph problems (which include densest subgraph and max-cut) by combining our edge-sampling techniques with the results of [EHW15]; see Section˜4.4 for details.
The min-cut problem illustrates a qualitative difference between the value oracle and comparison oracle settings: in the former, we can evaluate the expression $\nicefrac{{1}}{{2}}(w(\partial u)+w(\partial v)-w(\partial\{u,v\}))$ to learn whether edge $(u,v)$ exists (and what its weight is, if the graphs are weighted). In contrast, reconstructing a graph is not straightforward with cut comparisons. (E.g., all non-trivial cuts in $K_{3}$ have the same value, and the same holds for its complement graph $\bar{K}_{3}$, so we cannot distinguish them by comparison queries alone. Moreover, weighted graphs cannot be recovered at all using comparisons, even up to scaling (see Section˜C.4.2)! Given this, Theorem˜1.6 comes as a surprise.
The main idea behind both results is the following: for a vertex $v$ and set $S\not\ni v$, comparing $w(\partial(S+v))$ and $w(\partial(S))$ tells us whether more than half of $v$’s incident edges go into $S$. We build on this observation to identify all edges incident to $v$: for a sequence of nested sets $\emptyset\neq S_{0}\subseteq S_{1}\subseteq\ldots\subseteq S_{n-1}\neq V\setminus\{v\}$, we ask the above question for each $S_{i}$ to find a “tipping point”—a query where the sign of $w(\partial(S+v))-w(\partial(S))$ changes—whereupon we find an edge. We then refine the process to make edge recovery more efficient, and to perform efficient sampling on $G$ and get a randomized algorithm using only $\widetilde{O}(n)$ comparison queries; see Section˜4.
Weighted Min-Cut. These techniques do not extend to weighted graphs. E.g., there are many weighted graphs which cannot be identified using just comparison queries. (E.g., adding tiny-weight edges to a graph changes the graph but not the comparisons.) Also, while our techniques can give non-adaptive algorithms using $\operatorname{poly}(n)$ comparisons for min-cut/graph recovery in simple graphs, any non-adaptive algorithm for the weighted case must have exponential query complexity in the worst case. (See Section˜C.4 for a discussion of this and other bottlenecks.) That said, we give the following result for the case where the graph has a small number of distinct vertex degrees. (See Section˜C.3 for a proof.)
@@
Note that knowing the order of element weights is enough to optimize over matroids (using the greedy algorithm), but how can we compare element weights? We observe that if two elements share a circuit, then we can find two bases which differ on exactly these elements, and hence compare them. We then show that one can decompose any matroid so that we only need to compare such pairs of elements. (See Section˜D.1.)
For matroid intersection (and its special case, bipartite matchings), sorting the element weights is neither doable nor sufficient. Instead, we implement the shortest augmenting path algorithm (and its natural extension for matroid intersection) using only comparisons between matchings (i.e., common independent sets); see Section˜D.2. Finally, we show that the Bellman-Ford algorithm for shortest $s$-$t$ paths can be implemented using comparisons between $s$-$t$ walks to prove (iii).
Organization. In Section˜2, we present our first result on query complexity for Boolean Linear Optimization. Next, Section˜3 introduces our Global Subspace Learning framework, and gives efficient algorithms for specific classes of set systems. We then present efficient algorithms for other classic combinatorial problems: Section˜4 is dedicated to min-cuts, while Appendix˜D covers minimum-weight matroid bases, matroid intersections, and shortest $s$-$t$ paths. To improve flow, many technical proofs and discussions are deferred to the appendix.

### v2

@@ Abstract
    Having resolved the query complexity, we turn to the algorithmic question. We develop two general algorithmic frameworks:
    1. (a)
      Optimization from Certification: We present a novel Dual Ellipsoid framework that establishes an efficient reduction from optimization to certification. This framework demonstrates that to optimize efficiently, it is sufficient to design an efficient certification for the optimality of a candidate set $S^{*}$ with the knowledge of $w^{*}$ using only comparisons between feasible sets. Moreover, this framework also yields a deterministic low query complexity algorithm.
    2. (b)
      Global Subspace Learning (GSL): Tailored for integer objective functions bounded by $B$, we sort all feasible sets using only $O(nB\log(nB))$ queries, improving upon the $\widetilde{O}(n^{2})$ bound when $B=o(n)$. We efficiently implement this framework for linear matroids via algebraic techniques, yielding efficient algorithms with improved query complexity for problems like $k$-SUM, SUBSET-SUM, and $A+B$ sorting.
@@
    Our third set of results gives the first polynomial-time, low-query algorithms for several classic combinatorial problems. We develop such algorithms for finding minimum cuts in simple graphs, minimum weight spanning trees (and matroid bases in general), bipartite matching (and matroid intersection), and shortest $s$-$t$ paths.
Our work provides the first general query complexity bounds and efficient algorithmic results for this fundamental model.
Google Research. Emails: {vcohenad,gurug,renatoppl,jschnei}@google.comBocconi University. Email: tommasodorsi@gmail.comNew York University. Emails: {anupam.g,madhusudhan.p}@nyu.edu
§University of Michigan. Email: euiwoong@umich.eduDuke University. Email: debmalya@cs.duke.eduCarnegie Mellon University. Email: dwoodruf@andrew.cmu.edu
@@ 1.1.1 Query Complexity
1.1.2 Optimization from Certification
While the approach of Theorem˜1.1 provides strong query complexity bounds, we also provide a powerful alternative perspective via certification. In Section˜3, we show an efficient reduction from optimization to certification using a Dual Ellipsoid Algorithm. This framework reduces the optimization problem to designing a Conic-Certification Oracle (CCO). If one can efficiently certify the optimality of a candidate solution (with the knowledge of $w^{*}$) using only comparisons, one can efficiently optimize. Moreover, this framework also yields a deterministic algorithm that uses $\operatorname{poly}(n)$ queries for the Boolean Linear Optimization problem.
Theorem 1.3 (Optimization from Certification).
There is a deterministic algorithm that, given access to a Conic-Certification Oracle $\mathrm{(CCO)}$ for a family $\mathcal{F}$ represented by points $\mathcal{P}\subseteq\mathbb{R}^{d}$ with bit complexity $\langle\mathcal{P}\rangle$, finds the optimal solution $S^{*}=\arg\min_{S\in\mathcal{F}}\langle w^{*},\mathbbm{1}_{S}\rangle$ using $O(d^{3}\langle\mathcal{P}\rangle)$ calls to the $\mathrm{CCO}$ and separating hyperplane updates.
See Section˜3.5 for the proof of Theorem˜1.3. The fact that this framework also yields a deterministic algorithm that uses $\operatorname{poly}(n)$ queries for the BLO problem follows from ˜3.9. We demonstrate the power of this reduction by using it to efficiently find minimum weight perfect matchings in complete bipartite graphs.
@@ 1.1.3 Global Subspace Learning
Alongside the certification reduction, we develop a second, general algorithmic tool tailored for the case where the weights in the objective function are integers in the range [B,B][-B,B]. This Global Subspace Learning (GSL) algorithm is highly intuitive, and in fact, sorts all the feasible sets by their weight (in $\operatorname{poly}(|\mathcal{F}|,n)$ time for now):
@@
The main insight behind this result is the following: two sets $S,T\in\mathcal{F}$ with the same weight satisfy $\langle w^{*},\mathbbm{1}_{S}-\mathbbm{1}_{T}\rangle=0$, and hence the vector $(\mathbbm{1}_{S}-\mathbbm{1}_{T})$ gives us information about the subspace orthogonal to $w^{*}$. Our algorithm maintains a subspace $\mathcal{A}$ spanned by these orthogonal directions $(\mathbbm{1}_{S}-\mathbbm{1}_{T})$, and uses it to infer the weight class of other sets in $\mathcal{F}$ without making any additional comparisons; the details appear in Section˜4. While Theorem˜1.4 relies on weights being integral and bounded, it highlights how the structure of linear optimization can be exploited.
This explicit linear-algebraic structure allows us to get efficient algorithms from Theorem˜1.4:
@@
Moreover, we get new algorithms for the Subset Sum, $k$-Sum and Sorting $P+Q$ problems studied by [KLM19], with better bounds for the case when $B=o(n)$; see Section˜4.3 for details.
@@ 1.1.4 Efficient algorithms for Classical Combinatorial Families
Finally, we show that cut comparisons suffice to construct sparsifiers for a broad class of graph problems called heavy subgraph problems (which include densest subgraph and max-cut) by combining our edge-sampling techniques with the results of [EHW15]; see Section˜5.4 for details.
The min-cut problem illustrates a qualitative difference between the value oracle and comparison oracle settings: in the former, we can evaluate the expression $\nicefrac{{1}}{{2}}(w(\partial u)+w(\partial v)-w(\partial\{u,v\}))$ to learn whether edge $(u,v)$ exists (and what its weight is, if the graphs are weighted). In contrast, reconstructing a graph is not straightforward with cut comparisons. (E.g., all non-trivial cuts in $K_{3}$ have the same value, and the same holds for its complement graph $\bar{K}_{3}$, so we cannot distinguish them by comparison queries alone. Moreover, weighted graphs cannot be recovered at all using comparisons, even up to scaling (see Section˜D.4.2)! Given this, Theorem˜1.7 comes as a surprise.
The main idea behind both results is the following: for a vertex $v$ and set $S\not\ni v$, comparing $w(\partial(S+v))$ and $w(\partial(S))$ tells us whether more than half of $v$’s incident edges go into $S$. We build on this observation to identify all edges incident to $v$: for a sequence of nested sets $\emptyset\neq S_{0}\subseteq S_{1}\subseteq\ldots\subseteq S_{n-1}\neq V\setminus\{v\}$, we ask the above question for each $S_{i}$ to find a “tipping point”—a query where the sign of $w(\partial(S+v))-w(\partial(S))$ changes—whereupon we find an edge. We then refine the process to make edge recovery more efficient, and to perform efficient sampling on $G$ and get a randomized algorithm using only $\widetilde{O}(n)$ comparison queries; see Section˜5.
Weighted Min-Cut. These techniques do not extend to weighted graphs. E.g., there are many weighted graphs which cannot be identified using just comparison queries. (E.g., adding tiny-weight edges to a graph changes the graph but not the comparisons.) Also, while our techniques can give non-adaptive algorithms using $\operatorname{poly}(n)$ comparisons for min-cut/graph recovery in simple graphs, any non-adaptive algorithm for the weighted case must have exponential query complexity in the worst case. (See Section˜D.4 for a discussion of this and other bottlenecks.) That said, we give the following result for the case where the graph has a small number of distinct vertex degrees. (See Section˜D.3 for a proof.)
@@
Note that knowing the order of element weights is enough to optimize over matroids (using the greedy algorithm), but how can we compare element weights? We observe that if two elements share a circuit, then we can find two bases which differ on exactly these elements, and hence compare them. We then show that one can decompose any matroid so that we only need to compare such pairs of elements. (See Section˜E.1.)
For matroid intersection (and its special case, bipartite matchings), sorting the element weights is neither doable nor sufficient. Instead, we implement the shortest augmenting path algorithm (and its natural extension for matroid intersection) using only comparisons between matchings (i.e., common independent sets); see Section˜E.2. Finally, we show that the Bellman-Ford algorithm for shortest $s$-$t$ paths can be implemented using comparisons between $s$-$t$ walks to prove (iii).
Furthermore, by utilizing the Dual Ellipsoid reduction developed in Section˜3, we obtain an efficient algorithm for finding optimal perfect matchings in complete bipartite graphs.
Theorem 1.10 (Bipartite Perfect Matching).
There is a polynomial-time algorithm to find the minimum-weight perfect matching in a complete bipartite graph $K_{n,n}$ using only comparisons between perfect matchings.
Organization.
In Section˜2, we present our first result on query complexity for Boolean Linear Optimization. Next, Section˜3 presents our reduction from optimization to certification via the Dual Ellipsoid method. Section˜4 introduces our Global Subspace Learning framework, and gives efficient algorithms for specific classes of set systems. Section˜5 is dedicated to efficient algorithms for finding minimum cuts using cut comparisons. To improve flow, many technical proofs and discussions are deferred to the appendix, which includes the omitted proofs from the Dual Ellipsoid framework (see Appendix˜B), as well as our algorithms for minimum-weight matroid bases, matroid intersections, and shortest $s$-$t$ paths (see Appendix˜E).